OpenStudy (anonymous):

If the first equation is multiplied by 3 and then the equations are added, the result is _____. 3x + y = 3 x - 3y = -2 a. 10x = 11 b. 10x = 7 c. 8x = 7

OpenStudy (anonymous):

The problem is telling you what to do?

OpenStudy (anonymous):

Do you know how to use the elimination property?

OpenStudy (campbell_st):

here is the 1st equation multiplied by 3 9x + 3y = 9 + x -3y = -2 ------------------- just add them for the answer

OpenStudy (anonymous):

oh so is the answer 11 a.

OpenStudy (anonymous):

so its the first answer

OpenStudy (anonymous):

10x=7

OpenStudy (anonymous):

thank you! =))

OpenStudy (anonymous):

I was wrong, it is 10x = 7